When 1.00 kg lead specific heat =0.13jg °c at 100.0°c is adde to a quantity of water at 28.5, The final temperature of the lead water mixture is 35.2°c what is the mass of water present?


1
Expert's answer
2021-11-13T01:33:37-0500

Solution:

This problem can be summarized thusly:

qlost by lead = -qgained by water

 

q = m × C × ΔT

q = m × C × (Tf - Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C-1 g-1)

Tf = final temperature (°C)

Ti = initial temperature (°C)


Cwater = 4.186 J °C-1 g-1

Clead = 0.130 J °C-1 g-1


Therefore,

qlost by lead = (1000 g) × (0.130 J °C-1 g-1) × (35.2 - 100.0)°C = -8424 J

qlost by lead = -8424 J


qgained by water = mwater × (4.186 J °C-1 g-1) × (35.2 - 28.5)°C = mwater × 28.0462 J

qgained by water = mwater × 28.0462 J


qlost by lead = -qgained by water

-8424 = -28.0462 × mwater

mwater = (8424) / (28.0462) = 300.36 g = 300 g

mwater = 300 g


Answer: The mass of water is 300 grams